The sum of two numbers is 64. Twice the smaller number is 10 less than the larger number. Find the numbers.
Answer by algebrahouse.com(1659)

"The sum of two numbers is 64. Twice the smaller number is 10 less than the larger number. Find the numbers."

x = one number
64 - x = other number {sum of two numbers is 64}

2x = 64 - x - 10 {twice smaller is ten less than larger}
2x = 54 - x {combined like terms}
3x = 54 {added x to both sides}
x = 18 {divided both sides by 3}
64 - x = 46 {substituted 18, in for x, into 64 - x}

the two numbers are 18 and 46
www.algebrahouse.com